OpenStudy (anonymous):

What is the vertex of the function y=|-3x+2|-4

OpenStudy (triciaal):

are you sure the original is correct? the vertex is usually the highest or lowest point for a parabola, a function in the form ax^2 + bx + c

OpenStudy (triciaal):

possibility when x is negative y = 3x + 2 - 4 = 3x - 2 when x is positive y = -3x +2 - 4 = -3x -2 modulus usually means take the positive value y = -3x +2 -4 = -3x - 2 when x = 0 y = -2 when y = 0 then 4 = modulus -3 x + 2

OpenStudy (triciaal):